OpenStudy (anonymous):

Solve using the square root method. 9x2 - 25 = 0 A.{-5/3, 5/3} B.{-3/5, 3/5} C.{-3, 5} D.{-5, 3}

OpenStudy (anonymous):

option A

OpenStudy (eyust707):

Arnab, in the future spread the intended spirit of OpenStudy and comply with the Code of Coduct.

OpenStudy (anonymous):

9x^2 - 25 = 0 (3x-5)(3x+5) = 0 3x - 5 =0 3x+ 5= 0 3x = 5 3x= -5 x=5/3 x= -5/3 So, the answer is A.